A 1.9 m x 0.98 m x 5.7 cm thick solid wood door has a 0.4 m square window. The window has two panes of 0.45 cm thick glass with a 0.2 cm thick space between the glass panes that is filled with argon. Argon is used since it has a thermal conductivity that is 34% lower than that of air. The inside air temperature is 22
Answer #1

Solution:

The basic equation for heat conduction is Q = k.A \times \frac{\Delta T}{\Delta x}------1

Delta T is the temperature difference i.e = 22-(-10) = 32 celcius = 305.15 kelvin

Delta x is the thickness which the argon is filled = 0.2 cm = 0.002m

A is the area, From the problem,(1.9 x 0.98) = 1.862 meter.

The thermal conductivity of air approximately 0.025 w/m K. The conductivity of argon is 34% lesser therefore 0.085w/mK.

Substituting these values into equation one gives Q = 0.085\times 1.862 \times \frac{305.15}{0.002}

Q = 24148.04 watts is the heat flow through the argon gas.